OpenStudy (anonymous):

What is the product in simplest form? State any restrictions on the variable. y^2/y-3 x y^2-y-6/y^2+1y

OpenStudy (jack1):

throw some brackets in there, its a mess and can be interpreted too many different ways

OpenStudy (jack1):

also, there's an equation button just below where you type replier which makes it heaps easier to read and understand questions

OpenStudy (anonymous):

\[\frac{ y^2 }{ y-3 } \times \frac{ y^2 - y- 6 }{ y^2+1y}\]

OpenStudy (jack1):

ok \[\frac{ y(y+2) }{ y+1 }\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (jack1):

all good, understand how to get there?

OpenStudy (anonymous):

not exactly, could you explain?

OpenStudy (jack1):

top line (numerator) : y^2 (y^2 - y - 6)

OpenStudy (jack1):

becomes y^4 -y^3 -6y^2

OpenStudy (jack1):

bottom line (denominator): (y-3) (y^2 + y)

OpenStudy (jack1):

becomes y^3 +y^2 -3y^2 -3y =y^3 -2y^2 -3y

OpenStudy (anonymous):

That is very helpful, thank you for explaining it to me

OpenStudy (jack1):

all of the above works because when you multiply a fraction by a fraction, you times the numerators together and times the denominators together so our equation now reads:

OpenStudy (jack1):

all good, you understand now?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (jack1):

sweet as